A right triangle has a hypotenuse of length 15 and a leg of length 9. What is the length of the other leg of the triangle?
Answer by ReadingBoosters(3246)

Using,
a^2 + b^2 = c^2
a^2 + 9^2 = 15^2
a^2 + 81 = 225
a^2 + 81 - 81 = 225 - 81
a^2 = 144
sqrt(a^2) = sqrt(144)
a = 12